OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write the solution as an ordered pair. I did it but had a question about my answer

OpenStudy (anonymous):

m+5f=25 m+10f=40 subtract the first equation from the second 5f=15 f=3 m=10 (3, 10)

OpenStudy (anonymous):

is that correct? (:

OpenStudy (anonymous):

I didnt know which one would be x and which one would be y..... :P

OpenStudy (gw2011):

Your answer is correct. The session fee is $3 and the membership fee is $10. Use the letter m for membership fee and the letter f for the session fee. If for some reason you want to use the letters x and y to represent the different types of fees, you can do the following: Let x = membership fee Let y = session fee If you do this, you get: x + 5y = 25 x + 10y = 40 Now you can solve. You can also do this: Let x = session fee Let y = membership fee This will give you: 5x + y = 25 10x + y = 40 Now you can solve this. So, you can see that as long as you are consistent with what the letters represent and set up the equations correctly, you will get the same answers.